The Bob of a large pendulum has a mass of 30kg. The change in height of the pendulum as it swings is 0.5m.
A. Calculate the changes in GPE as the Bob moves from its highest to its lowest point.
B. All the GPE stored in the Bob at its highest point is transferred to kinetic energy as the Bob reaches its lowest point. Calculate the maximum speed of the bob.
1
Expert's answer
2019-11-12T17:26:41-0500

a.


"\Delta E_p=mgh=(30)(9.8)(0.5)=147\ J"

b.


"v=\sqrt{2gh}=\sqrt{2(9.8)(0.5)}=3.13\frac{m}{s}"